OpenStudy (anonymous):

Find the lim [(ln(x+5) - ln5) / x ] as x approaches 0

OpenStudy (anonymous):

looks like a l'hopital problem

OpenStudy (anonymous):

ok so there is an interesting rule that applies called L'hopital's rule. Since both top and bottom =0 if you plug in 0 for x, you can take the derivative of top and bottom until something doesn't equal zero. In this case, you would get \[d(\ln(x+5)-\ln(5)/x)=d(\ln(x+5)-\ln(5))/d(x)\] which is \[1/(x+5)/1|x=0\] which is \[1/5\]

OpenStudy (anonymous):

well actually not really. this is also the derivative of log at x = 5

OpenStudy (anonymous):

and since \[\frac{d}{dx}\ln(x)=\frac{1}{x}\] you get \[\frac{1}{5}\]

OpenStudy (anonymous):

I actually haven't learned how to find derivatives yet. Is there a way to do it without that?

OpenStudy (anonymous):

l'hopital works too of course. if you cannot use derivatives, then you can only guess

OpenStudy (anonymous):

no not really. you will have to plug in numbers close to 0 and see what you get

OpenStudy (anonymous):

take \[\frac{\ln(5.1)=\ln(5)}{.1}\] for example

OpenStudy (anonymous):

i meant \[\frac{\ln(5.1)-\ln(5)}{.1}\]

OpenStudy (anonymous):

without using derivatives you can neither use l'hopital not recognize this as a derivative you cannot do it with algebra

OpenStudy (anonymous):

*nor

OpenStudy (anonymous):

ok, gotcha, thanks(: I'l just do the plug in method.

OpenStudy (anonymous):

yup and see that you get numbers close and closer to 0.2